OpenStudy (anonymous):

how do you solve this equation: x+1/5=2x-2/6

OpenStudy (anonymous):

is it a proportion?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{ x+1 }{ 5 }=\frac{ 2x-2 }{ 6 }\]like this?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

thats a proportion

OpenStudy (imstuck):

It is a proportion, technically. It's also a ratio...technically. Solve it by cross-multiplication.

OpenStudy (anonymous):

are you sure?

OpenStudy (texaschic101):

positive

OpenStudy (anonymous):

well, that's a proportion. \[\frac{ a }{ b }=\frac{ c }{ d } \Rightarrow a \cdot d = b \cdot c\]

OpenStudy (anonymous):

Wonderful! Thank you guys!